OpenStudy (anonymous):

How can i solve (lnx)^2 = 1 ? I know that there is a way to move the logarithm to the other side of the equalty and it ends up like e^something, but i can't figure out the way to do that.

OpenStudy (lgbasallote):

first..take the square root of both sides \[\implies \sqrt{(\ln x)^2} = \sqrt 1\] simplify... \[\implies \ln x = \pm 1\] change to log form \[\implies \log_e x = \pm 1\] change to exponential form \[e^{\pm 1} = x\] therefore\[x = e ; \; x = \frac 1e\] does that help?

OpenStudy (phi):

e^(ln x)=x

OpenStudy (anonymous):

yeah, thank you lgbasallote that help me a lot :) I was actually stuck in the last step.

OpenStudy (lgbasallote):

you have been helped in the name of Jesus.

OpenStudy (lgbasallote):

^i always wanted to say that

OpenStudy (lgbasallote):

and welcome